OpenStudy (anonymous):

The slope-intercept equation of the line through the points (8,4) and (9,4) is y = mx+b what is M ? and what is B ?

OpenStudy (nowhereman):

Insert those two points as (x,y) and solve the resulting system of linear equations for m and b.

OpenStudy (anonymous):

im not sure how to get it, i got 4 for b , but how do i get m?

OpenStudy (nowhereman):

Solve the equation after you insert any point and 4 for b.

OpenStudy (anonymous):

you can get m by subtracting y1 from y2 over x1 from x2

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

you get b AFTER you get the slope

OpenStudy (amistre64):

ack!! that and it a straight line by the coordinates lol

OpenStudy (amistre64):

the slope of a horizontal line is 0... so the equation is: y = 0x +4 y = 4

OpenStudy (anonymous):

ok that makes sense, so m=0

OpenStudy (amistre64):

yes, m=0 :)

OpenStudy (anonymous):

ok so how would i do this one, (5.1,-2.3) and (-3.5,3.7) . whats the first thing i should do?

OpenStudy (amistre64):

first thing is to figure out the slope.. slope is defined by how steep the line rises. y1 -y2 slope = ------- x1 -x2

OpenStudy (amistre64):

-2.3 - (3.7) ---------- 5.1 -(-3.5)

OpenStudy (amistre64):

-6 --- is what I get if I added it right :) 8.6

OpenStudy (anonymous):

ok, yea thats right

OpenStudy (amistre64):

we can reduse that to make it easier to see: -6/8.6 = -60/86 = -30/43

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

now stick that slope into your equation for the line: y = mx + b y = (-30/43)x + b

OpenStudy (anonymous):

ok thanks that makes more sense

OpenStudy (amistre64):

pick any of the points they gave you and stick them in for x and y (x =-3.5, y=3.7) 3.7 = (-30/43)(-3.5) + b now solve for b

OpenStudy (anonymous):

ok thanks!!

OpenStudy (amistre64):

youre welcome.

OpenStudy (anonymous):

m is the slope and b is the y intercept, i.e the point lying on y axis